# 2-switch: transition and stability on forests and pseudoforests

### 08 Mar 2026 v1

@@ 1 Introduction
Let $G$ be a graph. We use $V(G)$ and $E(G)$ to refer to the vertex set and the edge set of $G$, respectively. We denote the order of $G$ by $|G|$ and the size of $G$ by $\lVert G\rVert$. The degree sequence of a graph $G$ with $V(G)=\{1,\ldots,n\}$ is denoted by $d=d(G)=(d_{1},\ldots,d_{n})$, where $d_{v}$ is the degree of the vertex $v$ in $G$. Let $a,b,c$ and $d$ be four distinct vertices of $G$ such that $ab,cd\in E$ and $ac,bd\notin E$. The process of deleting the edges $ab$ and $cd$ from $G$ and adding $ac$ and $bd$ to $G$ is referred to as a 2-switch on $G$. This is a classical operation, see [3, 7]. If $G^{\prime}$ is the graph obtained from $G$ by a 2-switch, it is straightforward to check that $d(G)=d(G^{\prime})$. In other words, this operation preserves the degree sequence. A key fact about 2-switch is the following.
Theorem 1.1.
@@
The realization graph of a degree sequence $d$ is the graph ${\cal G}(d)$ whose vertices are the graphs with degree sequence $d$, with two graphs $G,H$ forming an edge if $G$ and $H$ are obtained from one another by a $2$-switch (see [23]). A direct consequence of Theorem 1.1 is that $\mathcal{G}(d)$ is connected. In recent years, the problem of determining whether a given induced subgraph of $\mathcal{G}(d)$ is connected or not has been extensively studied. In this direction, many advances have already been made. Since Theorem 1.1 assures the existence of a path between $G$ and $H$ in $\mathcal{G}(d)$, a natural question to ask is how short that path can be. This is studied in [2] and more recently in [9]. Another natural question is whether there exists a version of Theorem 1.1 constrained to connected graphs. This was studied by Taylor in [18] and answered in the positive.
@@
One of the most studied problems in the literature is, given a graph parameter (clique number, domination number, matching number, etc.), finding the minimum and maximum values for the parameter on $V(\mathcal{G}(d))$ (see [6, 11, 12, 19, 21, 22]). Another interesting problem is deciding which values between the minimum and the maximum can be realized by a graph in $V(\mathcal{G}(d))$ (see [16, 17]). For instance, in [6] the authors study the matching number of trees with a given degree sequence, and of bipartite graphs with a given bipartite degree sequence. The authors find minimum and maximum values for the matching number in these families, and then show that all the other intermediate values are also attained. To establish the same property for bipartite degree sequences, they show that a 2-switch alters the matching number by at most $1$, and use the version of Theorem 1.1 for bipartite degree sequences. In the case of trees, the authors of [6] constructed the tree realizing each value between the minimum and the maximum, although they could have applied Theorem 1.2.
@@ 3 Forests
Let $F_{0}=F$ and $F_{i}=\tau_{i}(F_{i-1})$. As none of the vertices in $\Lambda$ is involved in any of the f-switches, $F_{i}-\Lambda=\tau_{i}(F_{i-1}-\Lambda)$ for all $i$. Hence $\theta$ is a sequence of 2-switches transforming $F$ into $F^{\prime}$. Since $F_{i}-\Lambda$ is obtained from $F_{i}$ by removing vertices of degree $1$, $F_{i}-\Lambda$ has the same cycles as $F_{i}$. Thus, as every $F_{i}-\Lambda$ is a forest, every $F_{i}$ is a forest as well. Therefore, $\theta$ is a sequence of f-switches transforming $F$ into $F^{\prime}$. ∎
@@
    Let $\ell$ be a leaf such that its neighbor $u$ in $F^{\prime}$ has degree at least $2$, and let $v$ be the neighbor of $\ell$ in $F$. If $\ell$ and $u$ are in different components of $F$, let $w$ be a neighbor of $u$ in $F$. If $\ell$ and $u$ are in the same component of $F$, consider the path $\ell v\ldots u$ in $F$. As $d_{u}\geq 2$, there is a neighbor $w$ of $u$ that is not in $\ell v\ldots u$. In either case, $\tau={{\ell\ v}\choose{u\ w}}$ is an f-switch on $F$ such that $\ell\in\Lambda(\tau(F),F^{\prime})$.
@@
        If every vertex in $F^{\prime}$ has degree $1$ choose a leaf $\ell\in V(F^{\prime})$. Else choose a leaf $\ell\in V(F^{\prime})$ such that its neighbor $u$ in $F^{\prime}$ has degree at least $2$.
@@
Note that in each step of (3) either $F=F^{\prime}$ or at least one leaf is removed (equivalently, one edge is removed). Thus, (3) runs at most $|E(F^{\prime})-E(F)|$ times. This number can be improved by $1$. First, we need a technical lemma.
@@
We proceed by contradiction. Assume that |E(G)E(H)|=1|E(G)-E(H)|=1. Clearly, $d(G)=d(H)$ implies $\lVert G\rVert=\lVert H\rVert$. Thus, |E(H)E(G)|=1|E(H)-E(G)|=1. Let $ab\in E(G)-E(H)$. As $d(G)=d(H)$, there must be an edge incident to $a$ and an edge incident to $b$ in E(H)E(G)E(H)-E(G). But, as |E(H)E(G)|=1|E(H)-E(G)|=1, this can only be possible if $ab\in E(H)$, contradicting the fact that $ab$ is the only edge in E(G)E(H)E(G)-E(H). ∎
@@
Let $U$ be a unicyclic graph and let $ab\in E(\operatorname{Cyc}(U))$, $cd\in E(\operatorname{For}(U))$. If $\tau$ is a 2-switch between $ab$ and $cd$, then there exists an edge $e\in\operatorname{Cyc}(U)$ such that $e\neq ab$ and $e\notin\tau(U-e)$.
@@
    Choose any $e\in E(\operatorname{Cyc}(U))$ and notice that GeG-e is a forest. Since $ab$ and $cd$ are in different components, $\tau(G-e)$ is a forest and so $\tau(G-e)+e$ contains at most one cycle. Since $|e\cap\{a,b,c,d\}|\leq 1$, $e\notin\tau(G-e)$. Hence, $\tau(G-e)+e=\tau(G)$.
@@
For each case of the hypothesis we have the following subcases: (A) $ab$ and $cd$ lie in the same component of $G$; (B) $ab$ and $cd$ lie in different components of $G$.
@@
For (1.B), let $ab\in E(U)$ and let $cd\in E(T)$, where $U$ is a unicyclic component and $T$ is a tree component. If $e\in E(U)-\{ab\}$, notice that $\tau$ is an f-switch on $(U-e)\dot{\cup}T$ and that $\tau(U\dot{\cup}T)=\tau((U-e)\dot{\cup}T)+e$ is a forest. Therefore, $\tau(G)$ is a pseudoforest. Finally, for (2.B), note that $\tau$ glues the two cycles containing $ab$ and $cd$ together in a new cycle. Hence, $c(\tau(G))=c(G)$. ∎
@@
    If $ab$ and $cd$ are in different components of $\operatorname{For}(U)$, for some unicyclic component $U$ of $G$, then $\tau$ is a p-switch if and only if it is a t-switch on UeU-e for all $e\in E(\operatorname{Cyc}(U))$.
@@ 7 Bipartite and non-bipartite graphs
Let $N_{k}$ be the non-bipartite graph with the following two components: a triangle $K_{3}$ together with a star $S_{k}$ of order $k$, for $k\geq 4$. Then, $d(N_{k})=(k-1)^{1}2^{3}1^{k-1}$. Observe that the only 2-switches we can perform on $N_{k}$ are between an edge of $K_{3}$ and an edge of $S_{k}$. Moreover, if $\tau$ is such a 2-switch, it is easy to see that $\tau(N_{k})$ is always a tree, that is, a bipartite graph. Thus, every 2-switch sequence transforming $N_{k}$ into a non-bipartite graph $N\not\approx N_{k}$ must pass through a bipartite graph. As an example of such an $N$, we can take the graph $N^{\prime}_{k}$ ($k\geq 4$), formed by 2 components: a path of order 3 and a triangle with k3k-3 leaves attached to one of its vertices (see Figure 2).
@@ 8 Stability and interval property
Let $\mathcal{X}$ be a connected induced subgraph of $\mathcal{G}(d)$ and let $\xi:V(\mathcal{X})\rightarrow\mathbb{Z}$. We say that $\xi$ is stable (under $2$-switch) in $\mathcal{X}$ if, for each $G\in V(\mathcal{X})$, we have
@@
Let $\mathcal{X}$ be a connected induced subgraph of $\mathcal{G}(d)$ and let $\xi:V(\mathcal{X})\rightarrow\mathbb{Z}$. Assume that one of the following inequalities holds for each graph $G\in V(\mathcal{X})$ and for every $2$-switch $\tau$ on $G$ such that $\tau(G)\in V(\mathcal{X})$:
@@ 8.2 Independence number and related parameters
Let $I$ be a maximum independent set in a graph $G$, $\tau={{a\ b}\choose{c\ d}}$ be a 2-switch on $G$ and let $V_{\tau}=\{a,b,c,d\}$. Notice that $|I\cap V_{\tau}|\leq 2$.
If $\left|I\cap V_{\tau}\right|\leq 1$, then $\alpha(\tau(G))\geq|I|=\alpha(G)$, because $I$ is an independent set in $\tau(G)$. We can easily conclude the same when $I\cap V_{\tau}=\{a,d\}$ or $I\cap V_{\tau}=\{b,c\}$.
If $I\cap V_{\tau}=\{a,c\}$, notice that $I$ is not an independent set in $\tau(G)$, since $ac\in E(\tau(G))$. Thus, IaI-a is an independent set in $\tau(G)$, and so $\alpha(\tau(G))\geq|I-a|=\alpha(G)-1$. The same argument holds if $I\cap V_{\tau}=\{b,d\}$.
@@ 8.3 Domination number
Assume $D$ is not a dominating set in $\tau(G)$. As the edges incident to vertices not in $\{a,b,c,d\}$ in $G$ and in $\tau(G)$ are the same, $D$ dominates every vertex in $V(G)-\{a,b,c,d\}$. Hence at least one vertex in $\{a,b,c,d\}$ is not dominated by $D$ in $\tau(G)$. Without loss of generality, assume that such a vertex is $a$ and consider its neighbors in $G$ and $\tau(G)$. Since the only edge incident to $a$ in $G$ that is not in $\tau(G)$ is $ab$, $b$ must be in $D$. Therefore, $d$ is dominated in $\tau(G)$ by $b$. Moreover, $D\cup a$ is a dominating set in $\tau(G)$, because $c$ is dominated by $a$ in $\tau(G)$. Thus, $\gamma(\tau(G))\leq|D\cup a|=\gamma(G)+1$.
@@ 8.5 Path-covering number and related parameters
    abab and $cd$ are both edges of the forest $\Phi$.
@@
    abab and $cd$ are not edges of the forest $\Phi$.
@@
    exactly one of $ab$ and $cd$ is an edge of the forest $\Phi$.

### 11 Mar 2026 v2

@@ 1 Introduction
Let $G$ be a graph. We use $V(G)$ and $E(G)$ to refer to the vertex set and the edge set of $G$, respectively. We denote the order of $G$ by $|G|$ and the size of $G$ by $\lVert G\rVert$. The degree sequence of a graph $G$ with $V(G)=\{1,\ldots,n\}$ is denoted by $d=d(G)=(d_{1},\ldots,d_{n})$, where $d_{v}$ is the degree of the vertex $v$ in $G$. Let $a,b,c$ and $d$ be four distinct vertices of $G$ such that $ab,cd\in E$ and $ac,bd\notin E$. The process of deleting the edges abab and cdcd from $G$ and adding acac and bdbd to $G$ is referred to as a 2-switch on $G$. This is a classical operation, see [3, 7]. If $G^{\prime}$ is the graph obtained from $G$ by a 2-switch, it is straightforward to check that $d(G)=d(G^{\prime})$. In other words, this operation preserves the degree sequence. A key fact about 2-switch is the following.
Theorem $1.$
@@
The realization graph of a degree sequence $d$ is the graph ${\cal G}(d)$ whose vertices are the graphs with degree sequence $d$, with two graphs $G,H$ forming an edge if $G$ and $H$ are obtained from one another by a 22-switch (see [23]). A direct consequence of Theorem 1.1 is that $\mathcal{G}(d)$ is connected. In recent years, the problem of determining whether a given induced subgraph of $\mathcal{G}(d)$ is connected or not has been extensively studied. In this direction, many advances have already been made. Since Theorem 1.1 assures the existence of a path between $G$ and $H$ in $\mathcal{G}(d)$, a natural question to ask is how short that path can be. This is studied in [2] and more recently in [9]. Another natural question is whether there exists a version of Theorem 1.1 constrained to connected graphs. This was studied by Taylor in [18] and answered in the positive.
@@
One of the most studied problems in the literature is, given a graph parameter (clique number, domination number, matching number, etc.), finding the minimum and maximum values for the parameter on $V(\mathcal{G}(d))$ (see [6, 11, 12, 19, 21, 22]). Another interesting problem is deciding which values between the minimum and the maximum can be realized by a graph in $V(\mathcal{G}(d))$ (see [16, 17]). For instance, in [6] the authors study the matching number of trees with a given degree sequence, and of bipartite graphs with a given bipartite degree sequence. The authors find minimum and maximum values for the matching number in these families, and then show that all the other intermediate values are also attained. To establish the same property for bipartite degree sequences, they show that a 2-switch alters the matching number by at most 11, and use the version of Theorem 1.1 for bipartite degree sequences. In the case of trees, the authors of [6] constructed the tree realizing each value between the minimum and the maximum, although they could have applied Theorem 1.2.
@@ 3 Forests
Let $F_{0}=F$ and $F_{i}=\tau_{i}(F_{i-1})$. As none of the vertices in $\Lambda$ is involved in any of the f-switches, $F_{i}-\Lambda=\tau_{i}(F_{i-1}-\Lambda)$ for all ii. Hence $\theta$ is a sequence of 2-switches transforming $F$ into $F^{\prime}$. Since $F_{i}-\Lambda$ is obtained from $F_{i}$ by removing vertices of degree 11, $F_{i}-\Lambda$ has the same cycles as $F_{i}$. Thus, as every $F_{i}-\Lambda$ is a forest, every $F_{i}$ is a forest as well. Therefore, $\theta$ is a sequence of f-switches transforming $F$ into $F^{\prime}$. ∎
@@
    Let $\ell$ be a leaf such that its neighbor $u$ in $F^{\prime}$ has degree at least 22, and let $v$ be the neighbor of $\ell$ in $F$. If $\ell$ and $u$ are in different components of $F$, let $w$ be a neighbor of $u$ in $F$. If $\ell$ and $u$ are in the same component of $F$, consider the path $\ell v\ldots u$ in $F$. As $d_{u}\geq 2$, there is a neighbor $w$ of $u$ that is not in $\ell v\ldots u$. In either case, $\tau={{\ell\ v}\choose{u\ w}}$ is an f-switch on $F$ such that $\ell\in\Lambda(\tau(F),F^{\prime})$.
@@
        If every vertex in $F^{\prime}$ has degree 11 choose a leaf $\ell\in V(F^{\prime})$. Else choose a leaf $\ell\in V(F^{\prime})$ such that its neighbor $u$ in $F^{\prime}$ has degree at least 22.
@@
Note that in each step of (3) either $F=F^{\prime}$ or at least one leaf is removed (equivalently, one edge is removed). Thus, (3) runs at most $|E(F^{\prime})-E(F)|$ times. This number can be improved by 11. First, we need a technical lemma.
@@
We proceed by contradiction. Assume that |E(G)E(H)|=1|E(G)-E(H)|=1. Clearly, $d(G)=d(H)$ implies $\lVert G\rVert=\lVert H\rVert$. Thus, |E(H)E(G)|=1|E(H)-E(G)|=1. Let $ab\in E(G)-E(H)$. As $d(G)=d(H)$, there must be an edge incident to $a$ and an edge incident to $b$ in E(H)E(G)E(H)-E(G). But, as |E(H)E(G)|=1|E(H)-E(G)|=1, this can only be possible if $ab\in E(H)$, contradicting the fact that abab is the only edge in E(G)E(H)E(G)-E(H). ∎
@@
Let $U$ be a unicyclic graph and let $ab\in E(\operatorname{Cyc}(U))$, $cd\in E(\operatorname{For}(U))$. If $\tau$ is a 2-switch between abab and cdcd, then there exists an edge $e\in\operatorname{Cyc}(U)$ such that $e\neq ab$ and $e\notin\tau(U-e)$.
@@
    Choose any $e\in E(\operatorname{Cyc}(U))$ and notice that GeG-e is a forest. Since abab and cdcd are in different components, $\tau(G-e)$ is a forest and so $\tau(G-e)+e$ contains at most one cycle. Since $|e\cap\{a,b,c,d\}|\leq 1$, $e\notin\tau(G-e)$. Hence, $\tau(G-e)+e=\tau(G)$.
@@
For each case of the hypothesis we have the following subcases: (A) abab and cdcd lie in the same component of $G$; (B) abab and cdcd lie in different components of $G$.
@@
For (1.B), let $ab\in E(U)$ and let $cd\in E(T)$, where $U$ is a unicyclic component and $T$ is a tree component. If $e\in E(U)-\{ab\}$, notice that $\tau$ is an f-switch on $(U-e)\dot{\cup}T$ and that $\tau(U\dot{\cup}T)=\tau((U-e)\dot{\cup}T)+e$ is a forest. Therefore, $\tau(G)$ is a pseudoforest. Finally, for (2.B), note that $\tau$ glues the two cycles containing abab and cdcd together in a new cycle. Hence, $c(\tau(G))=c(G)$. ∎
@@
    If abab and cdcd are in different components of $\operatorname{For}(U)$, for some unicyclic component $U$ of $G$, then $\tau$ is a p-switch if and only if it is a t-switch on UeU-e for all $e\in E(\operatorname{Cyc}(U))$.
@@ 7 Bipartite and non-bipartite graphs
Let $N_{k}$ be the non-bipartite graph with the following two components: a triangle $K_{3}$ together with a star $S_{k}$ of order $k$, for $k\geq 4$. Then, $d(N_{k})=(k-1)^{1}2^{3}1^{k-1}$. Observe that the only 2-switches we can perform on $N_{k}$ are between an edge of $K_{3}$ and an edge of $S_{k}$. Moreover, if $\tau$ is such a 2-switch, it is easy to see that $\tau(N_{k})$ is always a tree, that is, a bipartite graph. Thus, every 2-switch sequence transforming $N_{k}$ into a non-bipartite graph $N\not\approx N_{k}$ must pass through a bipartite graph. As an example of such an NN, we can take the graph $N^{\prime}_{k}$ ($k\geq 4$), formed by 2 components: a path of order 3 and a triangle with k3k-3 leaves attached to one of its vertices (see Figure 2).
@@ 8 Stability and interval property
Let $\mathcal{X}$ be a connected induced subgraph of $\mathcal{G}(d)$ and let $\xi:V(\mathcal{X})\rightarrow\mathbb{Z}$. We say that $\xi$ is stable (under 22-switch) in $\mathcal{X}$ if, for each $G\in V(\mathcal{X})$, we have
@@
Let $\mathcal{X}$ be a connected induced subgraph of $\mathcal{G}(d)$ and let $\xi:V(\mathcal{X})\rightarrow\mathbb{Z}$. Assume that one of the following inequalities holds for each graph $G\in V(\mathcal{X})$ and for every 22-switch $\tau$ on $G$ such that $\tau(G)\in V(\mathcal{X})$:
@@ 8.2 Independence number and related parameters
Let II be a maximum independent set in a graph $G$, $\tau={{a\ b}\choose{c\ d}}$ be a 2-switch on $G$ and let $V_{\tau}=\{a,b,c,d\}$. Notice that $|I\cap V_{\tau}|\leq 2$.
If $\left|I\cap V_{\tau}\right|\leq 1$, then $\alpha(\tau(G))\geq|I|=\alpha(G)$, because II is an independent set in $\tau(G)$. We can easily conclude the same when $I\cap V_{\tau}=\{a,d\}$ or $I\cap V_{\tau}=\{b,c\}$.
If $I\cap V_{\tau}=\{a,c\}$, notice that II is not an independent set in $\tau(G)$, since $ac\in E(\tau(G))$. Thus, IaI-a is an independent set in $\tau(G)$, and so $\alpha(\tau(G))\geq|I-a|=\alpha(G)-1$. The same argument holds if $I\cap V_{\tau}=\{b,d\}$.
@@ 8.3 Domination number
Assume $D$ is not a dominating set in $\tau(G)$. As the edges incident to vertices not in $\{a,b,c,d\}$ in $G$ and in $\tau(G)$ are the same, $D$ dominates every vertex in $V(G)-\{a,b,c,d\}$. Hence at least one vertex in $\{a,b,c,d\}$ is not dominated by $D$ in $\tau(G)$. Without loss of generality, assume that such a vertex is $a$ and consider its neighbors in $G$ and $\tau(G)$. Since the only edge incident to $a$ in $G$ that is not in $\tau(G)$ is abab, $b$ must be in $D$. Therefore, $d$ is dominated in $\tau(G)$ by $b$. Moreover, $D\cup a$ is a dominating set in $\tau(G)$, because $c$ is dominated by $a$ in $\tau(G)$. Thus, $\gamma(\tau(G))\leq|D\cup a|=\gamma(G)+1$.
@@ 8.5 Path-covering number and related parameters
    abab and cdcd are both edges of the forest $\Phi$.
@@
    abab and cdcd are not edges of the forest $\Phi$.
@@
    exactly one of abab and cdcd is an edge of the forest $\Phi$.